Starting in 2004, Donald and Bella have been purchasing Series EE bonds in their name to use for the higher education of their daughter Ashley, who currently is age 21. In 2019, they cash in $9,000 of the bonds to use for tuition, fees, and room and board. Of this amount, $3,000 represents interest. Of the $9,000, $8,100 is used for tuition and fees, and $900 is used for room and board. Donald and Bella's AGI, before the educational savings bond exclusion, is $131,250.

If an amount is zero, enter "0".

a. If Donald and Bella file a joint, how much is the savings bond exclusion?

Round any division to two decimal places and use rounded amount in subsequent computations. If required, round your final answer to the nearest dollar.

*Congress has attempted to assist low-to middle-income parents in saving for their children's college education. One of the ways Congress assists such families is through an interest income exclusion on educational savings bonds.*
Answer #1

The taxpayer can exclude the interest income if all conditions are satisfied.

  • The bonds were issued after 1989
  • Bond holder must be more than 24 years
  • The bonds were cashed during the tax year
  • Tax payer paid qualified higher education expenses
  • In case of jointly payment the AGI should be less than $153,500 and more than $123550

As all conditions are satisfied so interest income will be excluded.

Since, the bonds cash in for $9000($6000 principal and $3000 interest )

the amount of interest will reduced on pro rata basis

qualified educational expenses were 8100*

the interest exclusion will be taken = $8100/$9000=90%

Interest exclusion = 3000*90%=$2700

* Qualified educational expenses

only tution, fees will be taken no room and board will be considered.

So, now we will check the threshhold limit i.e. more than $123550 and less than $ 153500.

Since the AGI is $131250

The exclusion interest will be calculated below

THE amount exceed in AGI limit=$131250-123550

=$7700

The excess amount now divided with $30000 predetermined limit when the amount paid jointly.

7700/30000

=25.7%

=2700*25.7%

=693.9

So, the amount of interest excluded will be

2700-694= $2006

Amount of exclusion =$2006

And balance interest of 3000-2006=$994

Will be taxable.

Sorry for the inconvenience if you can provide me the correct value so,that i can check what's the issue.